(-x/(1-x)+(1+x)/x)/((1-x)/x-x/(1+x))
Answer by mananth(16949)

(-x/(1-x)+(1+x)/x)/((1-x)/x-x/(1+x)
LCM of numerator= x(1-x) LCM OF DENOMINATOR =x(1+x)
..
-x^2+(1+x)(1-x)/ x(1-x) / (1+x)(1-x)-x^2/x(1+x)
..
-x^2+(1+x)(1-x)/x(1-x) / x(1+x)/(1+x)(1-x) -x^2
..
-x^2+1-x^2 / x(1-x) / x(1+x)/ 1-x^2-x^2
..
1-2x^2 /x(1-x) / x(1+x)/ 1-2x^2
..
(1+x)/(1-x)